OpenStudy (anonymous):

inverse of g(x)=x-1/4?

OpenStudy (hysenberg):

find \[\frac{ 1 }{ g(x)}\]

OpenStudy (hysenberg):

g(x)=(4x-4)/4

OpenStudy (hysenberg):

now 1/g(x)=inverse of g(x)=4/(4x-4)=1/(x-1)

OpenStudy (anonymous):

so g-1(x)=4x-4

OpenStudy (hysenberg):

no g(x)=\[(4x-4)/4\]

OpenStudy (anonymous):

oh 4x+1

OpenStudy (hysenberg):

so \[g ^{-1}(x)=4/(4x-4)=1/(x-1)\]